OpenStudy (anonymous):

Misha used a probability simulator to pull 3 colored marbles from a bag and flip a coin 50 times. The results are shown in the tables below: Color of Marble Number of Times Rolled Blue 20 Green 18 Yellow 12 Heads Tails 28 22 Using Misha's simulation, what is the probability of pulling a blue marble and the coin landing tails up?

OpenStudy (anonymous):

Answers 410/2500 440/2500 32/50 and 42/50

OpenStudy (anonymous):

MEDAL!!!

OpenStudy (jtvatsim):

Here's the plan: 1. Figure out the probability of getting a blue marble based on the tables. 2. Figure out the probability of getting tails on the coin based on the tables. 3. To get the probability of blue marble AND tails on the coin, we multiply our answers to step 1 and 2.

OpenStudy (anonymous):

uh ok

OpenStudy (jtvatsim):

So, let's drill into the first part. "Probability of getting a blue marble." How many blue marbles did Misha get? Look at the table to find this.

OpenStudy (anonymous):

20

OpenStudy (jtvatsim):

Perfect! Next, we need to know how many marbles TOTAL there were.

OpenStudy (anonymous):

50

OpenStudy (jtvatsim):

That's right! So, the probability of getting a blue marble is found by taking: blue/total = 20/50 = 0.4 We just finished step 1.

OpenStudy (anonymous):

ok

OpenStudy (jtvatsim):

Now, for step 2, we basically do the same thing. How many tails were there? How many coin flips total?

OpenStudy (anonymous):

22

OpenStudy (jtvatsim):

Good on the tails. And if I'm counting correctly, I think there are 50 coin flips total.

OpenStudy (anonymous):

yes your right

OpenStudy (jtvatsim):

So that gives us tails/total = 22/50 = 0.44 Step 2 is done. Almost finished!

OpenStudy (anonymous):

xd ok

OpenStudy (jtvatsim):

Now, to get the probability of both happening (AND), we need to multiply the probabilities (if you are doing an OR problem, you would add the probabilities): So, we just have 0.4 * 0.44 = 0.176.

OpenStudy (jtvatsim):

Unfortunately, they were mean and gave us fraction answers, but all we do is stick them in the calculator and see which one is 0.176.

OpenStudy (anonymous):

ok

OpenStudy (jtvatsim):

And, with some number punching....... Aha! Looks like it is 440/2500 = 0.176.

OpenStudy (anonymous):

oh ok yes thank you so much

OpenStudy (jtvatsim):

You are welcome. Nice work!

OpenStudy (anonymous):

ty you too here a medal